# RaCalNet: Radar Calibration Network for Sparse-Supervised Metric Depth Estimation

### v1

@@ IV-A Datasets
1) ZJU-4DRadarCam Dataset [2]: We first evaluate our method on this public dataset, which was collected using a ground robotic platform equipped with an Oculii’s EAGLE 4D radar, a RealSense D455 camera, and a RoboSense M1 LiDAR sensor. It primarily covers diverse driving scenarios in both urban and off-road environments. The dataset contains 33,409 synchronized radar-camera keyframes, with 29,312 frames designated for training and validation, and 4,097 frames reserved for testing.
2) Real-World Deployment Scenarios: To further validate the generalizability of our approach beyond public datasets, we collected our own proprietary multimodal data in real-world deployment scenarios. These were captured using a ground robotic platform equipped with a ZF FRGen21 4D radar, a RealSense D435I camera, and a Livox HAP LiDAR which served as the source of ground truth depth labels (see Fig. 3). The data was collected primarily in campus environments and consists of 22,423 synchronized radar-camera keyframes, with 20,156 used for training and validation and 2,267 reserved for testing.
@@
Figure 4: Visualization of depth inaccuracies in the ZJU-4DRadarCam dataset. The top row illustrates misalignment in the LiDAR depth map (the original ground truth), primarily caused by calibration errors. The bottom row shows our recalibrated version of the ground truth, where the depth projection exhibits significantly improved accuracy.
IV-B Training Details and Evaluation Protocol
For the ZJU-4DRadarCam dataset, we addressed alignment inaccuracies stemming from imperfect sensor calibration (illustrated in Fig. 4) by optimizing extrinsic parameters and reprojecting the original LiDAR points. This recalibration process significantly improved geometric consistency between LiDAR measurements and RGB images. Throughout our experiments, our method exclusively employs sparse single-frame LiDAR supervision, while comparative methods maintain their original configurations using multi-frame projected and interpolated dense LiDAR for training. To validate methodological robustness in real-world deployment scenarios, we maintained identical supervision protocols: our approach utilized sparse single-frame LiDAR supervision, whereas all baselines employed their native training configurations relying on multi-frame densified LiDAR supervision.
@@ IV-B Training Details and Evaluation Protocol
For quantitative evaluation of depth estimation performance, we adopt standard metrics widely used in prior work [25], including Mean Absolute Error (MAE), Root Mean Square Error (RMSE), Absolute Relative Error (AbsRel), Squared Relative Error (SqRel), and thresholded accuracy ($\delta_{1}$).
@@ IV-C Evaluation on ZJU-4DRadarCam
We conducted a comprehensive evaluation of dense depth estimation performance across 50m, 70m, and 80m ranges. As reported in Table I, our RaCalNet framework consistently achieves state-of-the-art performance across all evaluated ranges, significantly outperforming the prior SOTA approach [2] using their best-performing configuration (DPT+Var+RC-Net). Specifically, RaCalNet achieves RMSE reductions of 35.08%, 29.12%, and 27.27% at 50m, 70m, and 80m ranges respectively. These improvements stem from our two-stage architecture: first identifying reliable radar points and refining pixel-level projections, followed by global optimization that aligns monocular predictions to metric scale while preserving structural integrity.
@@
Fig. 5 presents visual comparisons organized in five rows: (1) input image, (2) depth map predicted by Radarcam [2], (3) depth map predicted by our RaCalNet, (4) error map of Radarcam, and (5) error map of RaCalNet (all in meters). Red boxes highlight regions where RaCalNet exhibits noticeably better performance. The results show that our method accurately segments sky regions, delineates object boundaries with high precision, and effectively suppresses artifacts in complex scenes.
Figure 5: Experimental results for ZJU-4DRadarCam (from top to bottom): (a) Input image, (b) Radarcam’s [2] and our RaCalNet depth estimations, (c) Radarcam’s and our errors. All values are in meters (m). Red boxes highlight superior performance regions.
@@ IV-D Evaluation on Real-World Deployment Scenarios
Following the same evaluation protocol as in Section IV-C, we further validate the proposed framework using proprietary real-world deployment data. As summarized in Table II, our method consistently outperforms existing radar-guided depth estimation baselines [2] across all tested distance ranges.
TABLE II: Evaluations on Real-World Deployment (mm)
@@
TABLE III: Ablation of Our Module (mm)
@@
In particular, compared to [2], our method achieves substantial RMSE reductions of 46.24%, 43.42%, and 40.06% at 50m, 70m, and 80m intervals, respectively. These results set a new benchmark for long-range radar-based dense depth completion.
Figure 6: Experimental results for proprietary real-world data (from top to bottom): (a) Input image, (b) Radarcam’s [2] and our RaCalNet depth estimations, (c) Radarcam’s [2] and our errors. All values are in meters (m).
IV-E Ablation Study
We conduct detailed ablation studies on the real-world deployment data to analyze the contribution of individual components in our framework. Specifically, we evaluate three configurations: (1) the complete RaCalNet pipeline; (2) a variant retaining radar screening but without Pixel-wise Displacement Refinement; and (3) a variant excluding the entire Radar Refinement Module (both screening and displacement refinement). Results are shown in Table III and reveal the following insights:
1) Displacement Refinement Ablation: When retaining radar screening but disabling the pixel-level displacement refinement, we observe consistent performance drops (4.0-6.1% MAE increase across all ranges). This reveals two key insights: (i) even screened radar points require sub-pixel alignment adjustments due to residual calibration errors and temporal synchronization issues, and (ii) the learned displacement refinement effectively compensates for these imperfections while preserving the geometric validity of radar projections. The improvement is particularly notable at longer distances (70-80m) where calibration errors have more pronounced effects.
Figure 7: Qualitative comparison of 3D scene reconstruction results. (a) Input monocular RGB images. (b) Reconstructions based on depth estimated by our method. (c) Reconstructions using depth predicted by Radarcam. Our approach yields more complete and geometrically accurate results.
2) Impact of Radar Refinement: The complete removal of the Radar Refinement Module (both screening and displacement components) leads to severe performance degradation, with MAE increasing by 96.8% at 50m (2505mm vs. 1273mm) and 92.1% at 80m (2834mm vs. 1476mm). This demonstrates that raw radar measurements contain substantial noise and calibration errors that directly propagate to depth predictions. Our refinement module is crucial for: (i) identifying reliable radar points through cross-modal screening, and (ii) establishing accurate pixel-level correspondences, both being essential for maintaining metric accuracy in the fused output.
@@ IV-F Impact on Downstream Tasks: 3D Scene Reconstruction
In our experiment, we supply the same RGB images and ground-truth camera intrinsics across all methods to ensure fair comparison. Camera poses are left to be optimized by FrozenRecon during reconstruction. The only varying input is the predicted depth map, which is either produced by prior radar-guided depth estimation methods or by our full method integrating RaCalNet and the global optimization module.
@@ V CONCLUSIONS
We propose RaCalNet, a novel framework for dense metric depth estimation using radar and monocular images, which eliminates the need for dense LiDAR supervision. Instead, it leverages sparse LiDAR to supervise the refinement of raw radar points, generating accurate, pixel-wise depth anchors. These anchors are then used to calibrate scale and optimize monocular depth predictions, addressing key challenges such as radar unreliability and the distortion introduced by dense LiDAR interpolation.
Extensive experiments on public benchmarks and real-world scenarios demonstrate that RaCalNet significantly outperforms prior radar-guided methods, despite relying on only $\sim$1of the supervision density used in dense LiDAR-based approaches. The resulting depth maps preserve structural details and support downstream tasks such as 3D reconstruction. Our work introduces a new paradigm for radar-guided monocular depth estimation, providing a promising foundation for future research. In the future, we plan to explore temporal modeling across sequential frames to further improve robustness and accuracy.

### v2

@@ IV Experiments
To evaluate the effectiveness and transferability of RaCalNet, we conduct experiments on both a public dataset and a real-world deployment scenario. Our primary benchmark is the ZJU-4DRadarCam dataset [9], which offers time-synchronized 4D radar, RGB, and LiDAR data across diverse environments. In contrast, widely used datasets such as nuScenes [37] rely on 3D radar with only azimuth-range measurements on a fixed horizontal plane, lacking elevation information and thus unsuitable for evaluating 4D radar-based methods. To further assess real-world adaptability, we collect a proprietary dataset using a robotic platform equipped with production-grade sensors. Featuring distinct hardware configurations and capturing campus navigation scenarios, this real-world dataset allows for a more comprehensive evaluation of RaCalNet in practical applications.
For comparison, we adopt Radarcam [9] as our main baseline, as it represents a recent state-of-the-art method in radar-guided dense depth estimation. Radarcam leverages dense LiDAR supervision and multi-frame fusion, providing strong performance on their ZJU-4DRadarCam dataset. Its open-source implementation and methodological relevance to our approach make it a suitable and competitive reference point.
@@ IV-A Datasets
1) ZJU-4DRadarCam Dataset [9]: This public dataset was collected using a ground robotic platform equipped with an Oculii EAGLE 4D radar, a RealSense D455 camera, and a RoboSense M1 LiDAR. It includes 33,409 synchronized radar-camera keyframes, with 29,312 used for training and validation, and 4,097 reserved for testing. The dataset captures diverse scenes spanning both urban roads and off-road trails.
2) Real-World Deployment Scenarios: To further validate the generalizability of our approach, we collected a proprietary multimodal dataset using a different robotic platform equipped with a ZF FRGen21 4D radar, a RealSense D435I camera, and a Livox HAP LiDAR, which provides the ground-truth depth labels (see Fig. 3). Compared to the ZJU setup, this platform features different radar characteristics (e.g., field of view) and a LiDAR sensor with a unique scanning pattern. The dataset consists of 22,423 synchronized keyframes, primarily collected in campus navigation scenarios that emphasize dynamic pedestrians and varying lighting conditions. Among them, 20,156 frames are used for training and validation, and 2,267 for testing.
@@
IV-B Implementation Details and Evaluation Protocol
@@
For quantitative evaluation of depth estimation performance, we adopt standard metrics widely used in prior work [39], including Mean Absolute Error (MAE), Root Mean Square Error (RMSE), Absolute Relative Error (AbsRel), Squared Relative Error (SqRel), and thresholded accuracy ($\delta_{1}$). All metrics are computed separately within three evaluation ranges: 0-50 m, 0-70 m, and 0-80 m, to provide a comprehensive assessment across varying depths.
@@ IV-C Evaluation on ZJU-4DRadarCam
We initially evaluated our method and all baselines using the original ZJU-4DRadarCam dataset, strictly following the settings reported in Radarcam, where multi-frame densified LiDAR supervision is used. Under this configuration, our RaCalNet employed only sparse single-frame LiDAR supervision, while the comparative methods were trained with their default settings. Quantitative results under this setting are reported in Table I. While RaCalNet performed competitively, it did not exhibit significant improvements over the Radarcam (DPT+Var+RC-Net), which was contrary to our expectations given the design of our framework. Notably, our method produced visibly cleaner and more geometrically consistent depth maps, but these qualitative improvements were not fully captured by the reported metrics. This discrepancy prompted us to re-examine the ground-truth depth used for evaluation. We found that the dataset suffered from imperfect extrinsic calibration, resulting in noticeable misalignment between LiDAR point clouds and RGB images. Fig. 4 illustrates these geometric inconsistencies, which likely degraded the effectiveness of evaluation.
Figure 4: Visualization of depth inaccuracies in the ZJU-4DRadarCam dataset. The top row illustrates misalignment in the LiDAR depth map (the original ground truth). The bottom row shows our calibrated version of the ground truth, where the depth projection exhibits significantly improved accuracy.
@@
[…]
50m DORN [28] 2210.171 4129.691 0.157 939.348 0.783
Singh [11] 1785.391 3704.636 0.146 966.133 0.831
Radarcam [9] 1067.531 2817.362 0.087 575.838 0.922
Ours 1328.045 2248.606 0.095 340.544 0.893
70m DORN [28] 2402.180 4625.231 0.160 1021.805 0.777
Singh [11] 1932.690 4137.143 0.147 1014.454 0.828
Radarcam [9] 1157.014 3117.721 0.087 601.052 0.921
Ours 1492.442 2576.835 0.100 410.569 0.882
80m DORN [28] 2447.571 4760.016 0.161 1038.919 0.776
Singh [11] 1979.459 4309.314 0.147 1034.148 0.828
Radarcam [9] 1183.471 3228.999 0.088 610.501 0.920
Ours 1531.343 2659.084 0.101 428.687 0.880
To address this issue, we refined the sensor alignment by optimizing the extrinsic parameters and reprojecting the LiDAR data, which significantly improved the spatial consistency between LiDAR and camera views. We then re-trained both our method and the strongest-performing baseline on the corrected dataset using the same training configurations as before. Notably, RaCalNet continued to rely only on sparse single-frame LiDAR, while the baseline retained their original dense supervision.
After correction, overall performance increased, and RaCalNet achieved state-of-the-art results. As shown in Table II, our method achieved RMSE reductions of 35.08%, 29.12%, and 27.27% at 50m, 70m, and 80m respectively, compared to the strongest baseline. These results validate the effectiveness of both the correction process and our overall approach.
Figure 5: Experimental results for ZJU-4DRadarCam (from top to bottom): (a) Input image, (b) Radarcam’s and our RaCalNet depth estimations, (c) Radarcam’s and our errors. All values are in meters (m). Red boxes highlight superior performance regions.
TABLE II: EVALUATIONS ON ZJU-4DRADARCAM(CORRECTED) (mm)
@@
Figure 6: Experimental results for proprietary real-world data (from top to bottom): (a) Input image, (b) Radarcam’s and our RaCalNet depth estimations, (c) Radarcam’s and our errors. All values are in meters (m). Red boxes highlight superior performance regions.
Fig. 5 presents visual comparisons organized in five rows. Red boxes highlight regions where RaCalNet exhibits noticeably better performance. The results show that our method accurately segments sky regions, delineates object boundaries with high precision, and effectively suppresses artifacts in complex scenes.
The superior performance of RaCalNet can be attributed to its two-stage architecture. The Radar Recalibration module filters out unreliable radar points and refines their projected positions, reducing noise in the sparse depth prior. Subsequently, the Metric Depth Optimization module aligns monocular predictions with the corrected sparse ground truth, enforcing global metric consistency. Together, these components allow RaCalNet to produce depth maps that are not only accurate in scale but also structurally coherent.
@@ IV-D Evaluation on Real-World Deployment Scenarios
Following the same evaluation protocol as in Section IV-C, we further validate the proposed framework using proprietary real-world deployment data. As summarized in Table III, our method consistently outperforms Radarcam across all tested distance ranges. In particular, compared to Radarcam, our method achieves substantial RMSE reductions of 46.24%, 43.42%, and 40.06% at 50m, 70m, and 80m intervals, respectively. These results set a new benchmark for long-range radar-based dense depth completion.
TABLE III: EVALUATIONS ON REAL-WORLD DEPLOYMENT (mm)
@@
IV-E Ablation Studies
We conduct detailed ablation studies on the real-world deployment data to analyze the contribution of individual components in our framework. Specifically, we evaluate three configurations: (1) the complete RaCalNet pipeline; (2) a variant retaining radar screening but without Pixel-wise Displacement Refinement; and (3) a variant excluding the entire Radar Recalibration module (both screening and displacement refinement). Results are shown in Table IV and reveal the following insights:
TABLE IV: ABLATION OF OUR MODULE (mm)
@@ IV-E Ablation Studies
1) Displacement Refinement Ablation: When radar screening is retained but pixel-level displacement refinement is disabled, we observe consistent performance degradation, with MAE increasing by 4.0-6.1% across all depth ranges. This ablation highlights two key insights: (i) although unreliable radar points are removed through screening, the remaining points still suffer from slight misalignments caused by residual errors and temporal desynchronization between sensors, and (ii) the learned displacement refinement effectively compensates for these small yet impactful offsets, while preserving the geometric validity of radar projections.
2) Impact of Radar Refinement: The complete removal of the Radar Recalibration module (both screening and displacement components) leads to severe performance degradation, with MAE increasing by 96.8% at 50m (2505mm vs. 1273mm) and 92.1% at 80m (2834mm vs. 1476mm). This demonstrates that raw radar measurements contain substantial noise and errors that directly propagate to depth predictions. Our refinement module is crucial for: (i) identifying reliable radar points through cross-modal screening, and (ii) establishing accurate pixel-level correspondences, both being essential for maintaining metric accuracy in the fused output.
@@ IV-F Impact on Downstream Tasks: 3D Scene Reconstruction
In our experiment, we supply the same RGB images and ground-truth camera intrinsics across all methods to ensure fair comparison. Camera poses are left to be optimized by FrozenRecon during reconstruction. The only varying input is the predicted depth map, which is either generated by the baseline method Radarcam [9] or produced by our complete RaCalNet pipeline.
Figure 7: Qualitative comparison of 3D scene reconstruction results. (a) Input monocular RGB images. (b) Reconstructions using depth predicted by Radarcam, along with enlarged views of the highlighted regions. (c) Reconstructions based on depth estimated by our method, with corresponding enlarged views. Our approach yields more complete and geometrically accurate results.
@@ V Conclusions
We propose RaCalNet, a novel framework for dense depth estimation using radar and monocular images, which eliminates the need for dense LiDAR supervision. By leveraging sparse LiDAR to supervise the refinement of raw radar points, generating accurate, pixel-wise depth anchors. These anchors are then used to calibrate scale and optimize monocular depth predictions, addressing key challenges such as radar unreliability and the distortion introduced by dense LiDAR interpolation. Extensive experiments on public benchmarks and real-world scenarios demonstrate that RaCalNet achieves superior depth accuracy while using less than 1% of the supervision density required by dense LiDAR-based methods. The resulting depth maps not only preserve fine-grained structural details but also enhance downstream tasks such as 3D reconstruction.
While RaCalNet is currently not directly applicable to 3D radar datasets such as nuScenes [37] due to the lack of elevation information and the resulting planar point clouds, this limitation can be addressed through elevation inference and multi-view consistency. In future work, we also plan to incorporate temporal modeling to further enhance robustness across sequential frames.